Weekly Science Report 5: LMG02-05 Southern Ocean GLOBEC

Cruise days 26 to 31 August inclusive


BG-232-0 Costa/Burns/Crocker – Foraging Ecology of Crabeater seals

BG-234-0 Fraser – Winter Foraging Ecology of Adélie Penguins

BG-235-0 Fritsen – Sea Ice Microbial Communities

OG-241-0 Smith/Martinson/Perovich – Optical Environment of the Western Antarctic Peninsula Region

BG-244-0 Quetin/Ross – Winter Ecology of Larval Krill: Quantifying their interaction with the Pack Ice habitat


Cruise overview to date:


The week began with the LMG at Process Station #2 (Ice Station Rodin).  We had several more sightings of the elephant seal, which appeared late last week, including it hauling out just meters from the stern of the ship. A second, larger elephant seal was also seen.  Work on the ice station consisted of CTD casts in the morning and evening, scuba diving, ice coring and thickness transects, and the deployment of an ice drifter buoy.  Although several species of seals were seen in the area of open water aft of the ship during CTD casts, only the elephant seal hauled out.  Work at Process Site #2 was finished on 28 August and the LMG began to move in search of penguins to tag while awaiting the NBP's arrival in our area.  Unfortunately, no Adélies were seen, much less sampled.  On 31 August the two ships met up again and began to head northward toward Survey Station #4, which we hoped to make into Process Station #3. We conducted a transfer of personnel and equipment between ships and continued on our way.  Upon arrival at the northern edge of the grid we discovered that the storm from the previous few days had broken the large floes into smaller pieces ~10 meters in size.  In addition, we felt a significant swell propagating from the open ocean.  After much discussion about where to set up, it was decided that the LMG would move about on our own in this general area and pick floes to work for the day, rather than maintain an extended stay on any single floe.  This “free ship” mode of occupying a process station offers advantages as well as disadvantages, and it remains to be seen how productive it will be.  Our first day here has yielded success in finding krill.


Daily Activities:


26 August 2002

We continued to occupy Process station #2 (Rodin).  The weather changed considerably with 15-20 knot winds coming from the W to NW.  It was warmer with temperatures around -9°C.  A crabeater seal was spotted hauled out 400 m to our stern and several were seen in the open water immediately astern of the ship. Given the change in weather coupled with unstable ice conditions no attempt was made to capture the seal.  During the afternoon an elephant seal hauled out just aft of the ship and during the evening CTD two elephant seals were spotted in the open area immediately to the stern.  One could have been the adult female that hauled out, however, the 2nd individual was definitely an adult male southern elephant seal. Ice coring and optics measurements as well as scuba diving operations continued. The dive team, though collecting data from their underwater transects, still has not been able to find krill in direct association with the over-rafted ice.  They placed one-meter ring drift nets in the water overnight to attempt to find krill for growth experiments. They temporarily discontinued their diving operations today when a leopard seal was spotted in the open lead to the stern of the ship


27 August 2002

The LMG continued to occupy Process Station #2. The weather warmed up considerably with temperatures reaching a balmy -1 to -5°C.  Meanwhile, the Ice Coring team and the Krill team continue to collect data and samples.  We discussed with the NBP when and where to go next as work at this location has begun to wind down after 6 days here.  The CTD has been plagued with various arcane problems but we are still attempting to conduct two casts per day. The CTD casts seem to be a time when seals appear behind the stern of the vessel.  We have sighted several crabeater seals, a leopard seal, a Weddell seal and at least two elephant seals when we conducted CTD casts. Meanwhile, the Ice Coring team and the Krill team continue to collect data and samples. 

28 August 2002

The LMG completed work at Process Station #2 around noon today.  We pulled all our gear off the ice, except for the ice drifter buoy, and made our way toward a lead of open water/thin ice nearby.  The plan was to search for penguins to sample and tag, and to conduct a CTD cast.  Activities were coordinated with the NBP, but we were able to move somewhat independently.


29 August 2002

Today as the LMG pulled away from Process Station #2 or ice station “Rodin”, the ice team frantically scurried about the ice, collecting the last cores and thickness measurements from the piece of ice that we had called home for the last week. Other groups prepared for getting underway in search of penguins and continued analysis of samples already collected. After a final CTD cast we went cruising around the northern end of the GLOBEC grid in search of penguins. Although we did not have any success finding birds, we did find that we could move on our own.  Nevertheless, as it is unlikely that we could get to the next station on our own, we awaited the arrival of the NBP to assist in moving us to our third and final process station located at or near Survey Station #4.


30 August 2002

Today started out with the search for Penguins, however, by mid-morning the wind had picked up and the visibility dropped such that we could see only a few feet in front of the ship.  Since noon time the ship has been parked in the ice.  Parked, but still moving, as the ice pack has been drifting at 1.3 knots (nautical miles per hour), blown by 40 to 50 knot winds all afternoon. Our plan was to meet up with the NBP and proceed to our third Process station after transferring some gear and personnel between ships but this was delayed due to inclement weather.


31 August 2002  

The LMG and NBP met up in the morning and conducted a transfer of personnel and gear, this time with a zodiac.  The weather improved from the blizzard conditions and zero visibility of yesterday.  The penguin team was finally captured and tagged a bird.  Both ships proceeded to convoy to Survey Station #4 which we hope will serve as Process Site #3.


1 September 2002 

After reaching Survey Station #4, the LMG and NBP carried out a simultaneous CTD cast to allow for intercomparison of the various sensors on board the rosette. The LMG then made it on our own through the loose and recently broken up pack toward the southeast.  Just before lunch an acceptable floe was located to work on, and after a quick meal the ice team deployed to collect samples.  This work was followed by a scuba dive, which was quite successful in obtaining krill samples. The ice has been heavily fractured due to the heavy swell associated with the recent storm.  It is likely that will we go into a modified process station, where we sample different individual small floes each day.


Individual group reports: 


BG-232 – Seal Foraging Ecology- (Costa, Shaffer, Barnes, McDonald, and Kuhn)

Due to the cold weather very few if any seals have been sighted hauled out and none were sighted at the Ice Station.  A crabeater seal was spotted hauled out 400 m to our stern and several were seen in the open water immediately astern of the ship. However, given the distance to the seal and poor ice coupled with a weather report for high winds and low visibility no attempt was made to capture it.  Somewhat later in the afternoon an adult female elephant seal hauled out just aft of the ship (66° 23.7′S, 71° 07.5′W; 543 m).  We went out to take photographs and measure her using photogrammetry.   Although, we were able to get within 50 m of her, it appeared that a closer approach would likely send her into the water so we were unable to get a photo with a calibrated object nearby.  After the seal left in the late afternoon, Brett Pickering, Heidi Geisz (BG-234), Scott Shaffer, and Julie Barnes went to the site where the elephant seal had been hauled out.  They collected what appeared to be a vomit sample.  A number of unique parasites were observed, as well some fish otoliths. CTD casts seem to be a time when seals appear behind the stern of the vessel.  We have seen several crabeater seals and a leopard seal when we have conducted CTD casts and during the evening CTD two elephant seals were spotted in the open area immediately to the stern (66º 23.1′S, 70º 56.5′W; 549 m).  One could have been the adult female that hauled out earlier that day; however, the 2nd individual was definitely an adult male southern elephant seal.


We have completed analysis of all samples that can be done on board ship.  This includes blood hemoglobin, blood volume and blood count measurements. So far the animals tagged on this cruise have remained in the general area where they were tagged.  Unfortunately, now 4 of the 11 tags that were deployed have ceased transmitting.  We suspect that this is due to the heavy ice conditions breaking the antennae.  We are still planning to capture seals for stomach sampling as well as to collect additional body composition data.  In the meantime we are analyzing our data and helping the other science groups with ice coring, buoy deployments and diver tending.


BG-234 – Penguin Foraging Ecology (Heidi Geisz, Brett Pickering)

The second ice station remained unproductive for the penguin team.  No Adélies were spotted during the week.  To keep active team members assisted with dive and ice coring operations.  After leaving the second ice station without seeing a penguin, the Sea Bird group aboard the NBP invited our team to transfer to NBP for the duration of the upcoming third ice station.  It was decided that the probability of seeing Adélie Penguins aboard the NBP along its survey grid was higher than remaining on another ice station.  As luck would have it, soon after the transfer of personnel to the NBP was made a group of 30 penguins were spotted.  These birds scattered off the ice and into the open water, with a couple popping out on small floes.  One of these Adélies, a female, was captured and a dive depth transmitter was placed on her back.  


BG-235 - Sea Ice Microbial Communities (Adkins, Blees, Cunningham, Hartsough, Fritsen)

Ice collections and experiments continued throughout the remainder of Ice Station Rodin.  Specifically, over twenty ice cores were collected in efforts to determine integrated biomass of sea ice microalgal communities.  Several cores (ca. 4) were collected at the optics sites where the contribution of algal and particulate matter to the attenuation of visible irradiance through the ice and snow is being determined in conjunction with OG-241.


We have continued to find enhanced algal biomass (>20 micro grams per liter) in cores measuring over 70 centimeters in length.  Cores collected in the flat areas (with thicknesses less than 50 cm) have yielded a lower fraction of the cores having high biomass.  The distributions of the peaks in biomass in all cores have been found at all depths in ice core profiles.  Nutrient to salinity ratios are preliminary and show a slight depletion/non-conservative behavior that would be indicative of in situ production or ice formation in a more nutrient deplete water body.  Production assays on these communities indicate high potential for in situ production and in situ photoadaptation.  Hence, the early onset of the spring bloom may be occurring in the ice, which could be attributed in part to the relatively warm (average in situ temperatures over -3ºC) ice sheet with moderate in situ brine salinities (less than 60 psu in brine pockets).  Such conditions are not prohibitive for algal photosynthesis and growth.  Further, analysis and experimentation this week as we move to the northernmost ice stations could provide more conclusive evidence for the start of the spring bloom.


OG-241 Optical Environment of the Western Antarctic Peninsula Region (Elder and Claffey)

In the last week we have concluded ice operations at Ice Station “Rodin”. This included two optics stations, ice drifter deployment, 100 meter thickness transect with half of it re-measured 2 additional times to determine growth/ablation over the life of the station, and ice coring. Ice structure was performed on cores for the entire cruise thus far in the freezer van while in transit between stations.  Much of our time at ice station Rodin was spent at ­20°C temperatures, the ice however did not see this cold temperature and did not grow as much as expected.  This was also noted by the dive team (BG-244) as many of their under ice measurements indicated ablation. While in transit between locations, hourly ice observations were performed.  The ice near grid station #4 is very loose and recently broken. The breaks in the ice are fresh and un-eroded. There remains a swell traveling through the pack ice. This was once an area of massive floes, now, however, it is a conglomeration of small 15-20 meter pieces of floes that continually rub together in the swell. The station that we measured on 1 September that seemed highly ridges was in fact, relatively flat with massive snowdrifts.  Ice 50 cm thick with 30 cm drifts of snow. Being near the freezing point, the surface was flooded with 10 cm of water on top of the ice due to the weight of the snow added during the storm 2 days ago. Due to the ice conditions, we will be sampling ice floes in the immediate area but plans of revisiting the same ice floe more than once is not a viable option.


BG-244 - Winter Ecology of Larval Krill: Quantifying their interaction with the Pack Ice habitat (Quetin, Hessell, Willis, Boch, Oakes, Johnston and Dovel)

This week we finished work at Process Station Rodin (P.S.#2) and spent the remainder of the week in transit to the last and most northern Process Station.  At P.S. #2 we spent most of our dive time collecting suction samples from under-ice surfaces.  We collected approximately 50 random samples during the week for microscopic, chn, hplc and chl a analysis.  Some of the collections were coordinated with the topside ice collection efforts of BG-235.  Collecting these samples involved many people outside our project.  Thank you.  We could not have accomplished so much in the cold temperatures without additional help.  We hope to repeat this process at the northern station.  Krill were also collected, but in a more unusual way than on past cruises when we were able to collect larvae from ice surfaces with aquarium nets.  At P.S. #2 krill continued not to be associated directly with under-ice surfaces.  In an attempt to collect larvae and assess the plankton community within meters of the ice, we fished 1-m nets through our dive holes in the ice floe at 1 m and 10 m day and night.  We were successful at catching krill larvae, and did 5 day/night combinations of tows.  Based on observation of the catches, we caught many more larvae at night, and more larvae at a depth of 1 m compared to 10 m.  The krill catches allowed us to continue our experiments with growth and preserve samples for stage and length frequency and chn analysis.  Thanks again to all those who helped us this week.



Submitted by:

Daniel Costa, Chief Scientist and

Karl Newyear, Marine Projects Coordinator

LMG 02-05